# Radiative corrections to $\boldsymbol{\tau\to\pi\pi\nu_{\tau}}$

### v1

@@ 1 Introduction
In this work, we consider various aspects of the radiative corrections to $\tau\to\pi\pi\nu_{\tau}$ decays Colangelo et al. (2025), addressing a key challenge in the $\tau$-based HVP evaluation. After reviewing some basic formalism in Sec. 2, we recall the results of earlier calculations in chiral perturbation theory (ChPT) Cirigliano et al. (2001, 2002), whose validity we aim to extend beyond the low-energy region via a dispersive analysis, see Secs. 3 and 4, motivated by previous findings in the context of $e^{+}e^{-}\to\pi^{+}\pi^{-}$ Ignatov and Lee (2022); Colangelo et al. (2022b) that the corresponding structure-dependent virtual corrections can indeed be sizable. The matching to ChPT, as described in Sec. 5, also allows us to establish a suitable basis for the connection to short-distance (SD) contributions and lattice QCD Bruno et al. (2018), most conveniently expressed in terms of the chiral low-energy constants (LECs). For the real-emission diagrams, we critically revisit previous calculations using resonance models Flores-Baéz et al. (2006, 2007); Davier et al. (2010); Miranda and Roig (2020); Davier et al. (2024); Castro et al. (2025), both regarding the numerical implementation near threshold and the estimate of the resulting uncertainties, as described in Sec. 6. For the numerical evaluation, we then extend the dispersive representation of the pion form factor to include the $\rho^{\prime}$, $\rho^{\prime\prime}$ resonances, based on which we perform fits to the available data sets for the $\tau\to\pi\pi\nu_{\tau}$ spectral function from Belle Fujikawa et al. (2008), ALEPH Schael et al. (2005); Davier et al. (2014), CLEO Anderson et al. (2000), and OPAL Ackerstaff et al. (1999) in Sec. 7. The consequences for a $\tau$-based evaluation of the $2\pi$ contribution, $a_{\mu}^{\text{HVP, LO}}[\pi\pi,\tau]$, are discussed in Sec. 8, before concluding in Sec. 9. Further details on the radiative process $\tau\to\pi\pi\nu_{\tau}\gamma$ are relegated to App. A.
@@ 2 Formalism
Here and throughout we follow the convention that a pion without index is identified with a charged pion, in particular, $M_{\pi}\equiv M_{\pi^{\pm}}$ and thus $\beta_{\pi\pi}(s)=\sigma_{\pi}(s)$. The final factor in Eq. (12) refers to IB in the matrix elements, i.e., the pion VFF as measured in $e^{+}e^{-}$ scattering, $F_{\pi}^{V}(s)$, or the weak decay, $f_{+}(s)$. In the latter case, in principle, a second form factor $\tilde{f}_{-}(s,t)$ contributes to the matrix element, see Eq. (23) below, but only at second order in IB. At LO in IB, the $\tau$ decay rate can therefore be expressed as
@@
$\Delta(s,t)=1+2\tilde{f}_{+}(s,t)+g_{\text{rad}}(s,t)\,,\qquad g_{\text{rad}}(s,t)=g_{\text{Low}}(s,t)+g_{\text{rest}}(s,t)\,$, (18)
contains both virtual and real photon corrections, $\tilde{f}_{+}(s,t)$ and $g_{\text{rad}}(s,t)$, respectively, where the latter is further separated into the leading contribution from Low’s theorem Low (1958) and the rest.
@@ 3.1 General considerations
$\mathcal{M}=G_{F}V_{ud}^{*}f_{+}(s)\bar{u}(l_{2},\nu_{\tau})\gamma^{\mu}(1-\gamma_{5})u(l_{1},\tau)\left[(q_{2}-q_{1})_{\mu}\tilde{f}_{+}(s,t)+(q_{1}+q_{2})_{\mu}\tilde{f}_{-}(s,t)\right]\,$, (23)
where $\tilde{f}_{+}(s,t)$ represents the contribution due to the $J^{P}=1^{-}$ component of the weak current and $\tilde{f}_{-}(s,t)$ vanishes in the isospin limit. We have factored out $f_{+}(s)$, which is defined in analogy to Eq. (6)
@@
normalized in such a way that $f_{+}(s)=F_{\pi}^{V}(s)$ in the isospin limit. In these conventions, $\tilde{f}_{+}(s,t)\to 1$ in the limit in which the matrix element (23) is mediated by a single current insertion. In principle, Eq. (24) also involves a second form factor $f_{-}(s)$, which is related to $\tilde{f}_{-}(s,t)$ in the same single-current approximation and constitutes an example of a second-class current Weinberg (1958); Descotes-Genon and Moussallam (2014). As shown in Ref. Descotes-Genon and Moussallam (2014), $f_{-}(s)$ can be traded for a scalar form factor $f_{0}(s)$, with $f_{0}(0)=1$ and $f_{-}(s)=\frac{\Delta_{\pi}}{s}\big[f_{+}(s)-f_{0}(s)\big]$, which shows that $f_{-}(s)$ scales with the pion mass difference $\Delta_{\pi}=M_{\pi}^{2}-M_{\pi^{0}}^{2}$.
@@
$\tilde{f}_{+}^{\text{tree}}(s,t)=1\,,\qquad\tilde{f}_{-}^{\text{tree}}(s,t)=0\,$, (26)
since also $f_{+}(s)=1$ at LO in ChPT. The spin-averaged squared amplitude can be expressed by
$\frac{1}{2}\sum\limits_{\text{spins}}|\mathcal{M}|^{2}$ $=2G_{F}^{2}|V_{ud}|^{2}|f_{+}(s)|^{2}\Big\{m_{\tau}^{2}(m_{\tau}^{2}-s)|\tilde{f}_{-}(s,t)|^{2}$
$-2m_{\tau}^{2}(2M_{\pi^{0}}^{2}+m_{\tau}^{2}-s-2t)\text{Re}\,\left[\tilde{f}_{+}(s,t)\tilde{f}_{-}^{*}(s,t)\right]$
$+[4M_{\pi}^{2}(M_{\pi^{0}}^{2}-t)+4t(s+t-M_{\pi^{0}}^{2})-m_{\tau}^{2}(s+4t)+m_{\tau}^{4}]|\tilde{f}_{+}(s,t)|^{2}\Big\}\,.$ (27)
Furthermore, the differential decay width is given by
@@
with the phase space in the $\pi^{0}\nu_{\tau}$ invariant mass squared limited by $t_{\text{min}}(s)\leq t\leq t_{\text{max}}(s)$, see Eq. (16), and in the $\pi^{-}\pi^{0}$ invariant mass squared by $(M_{\pi}+M_{\pi^{0}})^{2}\leq s\leq m_{\tau}^{2}$. Reducing Eq. (23) to the single-current-insertion limit $f_{+}(s)\tilde{f}_{\pm}(s,t)\to f_{\pm}(s)$, see Eq. (24), the integral over $t$ reproduces Cirigliano et al. (2001)
@@ 3.2 Radiative corrections
At $\mathcal{O}(e^{2}p^{2})$ in ChPT, the diagrams arising from the LO effective Lagrangian of Ref. Knecht et al. (2000) are shown in Fig. 1. In this work, we employ dimensional regularization to deal with both the UV and the IR divergences arising in the computation. Following the definition in Eq. (23), the form factors $\tilde{f}_{+}$ for the diagrams in Fig. 1 can be expressed as
@@ 4 Dispersive analysis
As a first modification beyond the pure ChPT analysis summarized in Sec. 3, it is clear that the strong final-state interactions between $\pi^{-}$ and $\pi^{0}$ need to be taken into account, after all, the main purpose of studying this decay concerns a precision measurement of the corresponding pion VFF. This leads to a nontrivial $f_{+}(s)$ in Eq. (23), while in the ChPT analysis of Sec. 3 one has $f_{+}(s)=1$ at this order in the expansion.
@@ 4.2 Dispersive calculation of the box diagram
When supplementing the box diagram with unsubtracted dispersive representations of the pion VFF, see Eq. (34), and projecting it onto the $\tilde{f}_{+}(s,t)$ form factor, see Eq. (23), the following expression can be obtained
@@ 4.3 Numerical treatment of $\boldsymbol{D_{0}}$ functions
The dependence on the (high-energy) integral cutoff $\Lambda^{2}$ of $I_{\ell 1}$ and the one in the integral $\int\text{d}s^{\prime\prime}D_{0}^{\text{rest}}(t,s^{\prime\prime})$ in Eq. (54) cancels against the $\Lambda$ dependence introduced via $\theta(\Lambda^{2}-s^{\prime\prime})$ in the second term proportional to $p_{2}(s,t)$. The imaginary part that emerges from the Cauchy propagator, see Eq. (55), ultimately does not play a role, as the final result is determined by the interference with the real tree-level amplitude.
@@ 7 Fits to the $\boldsymbol{\tau}$ spectral function
$\Delta(s,t)=1+2\tilde{f}_{+}^{\text{match}}(s,t)+g_{\text{Low}}(s,t)+g_{\text{rest}}(s,t)\,.$ (112)
@@ 9 Summary and conclusions
A further main outcome of this work is the substantial reduction of the uncertainty in the $G_{\text{EM}}(s)$ contribution, leaving the matching between the short-distance factor $S_{\text{EW}}^{\pi\pi}$ and the radiative corrections described by $G_{\text{EM}}(s)$ as the dominant source of uncertainty in the $\tau$-specific IB corrections. This matching can be further improved using input from lattice QCD, and establishing the latter connection could also help address the remaining, most critical IB correction in the matrix elements. That is, our work allows for a reliable calculation of the long-range radiative corrections, but lattice-QCD techniques as well as complementary dispersive calculations of IB in the pion VFF are also needed to achieve a complete account of IB corrections to hadronic $\tau$ decays. Once all these ingredients become available, a robust evaluation of the two-pion HVP contribution to the anomalous magnetic moment of the muon will be possible.

### v2

@@ 1 Introduction
In this work, we consider various aspects of the radiative corrections to $\tau\to\pi\pi\nu_{\tau}$ decays Colangelo et al. (2025a), addressing a key challenge in the $\tau$-based HVP evaluation. After reviewing some basic formalism in Sec. 2, we recall the results of earlier calculations in chiral perturbation theory (ChPT) Cirigliano et al. (2001, 2002), whose validity we aim to extend beyond the low-energy region via a dispersive analysis, see Secs. 3 and 4, motivated by previous findings in the context of $e^{+}e^{-}\to\pi^{+}\pi^{-}$ Ignatov and Lee (2022); Colangelo et al. (2022b) that the corresponding structure-dependent virtual corrections can indeed be sizable. The matching to ChPT, as described in Sec. 5, also allows us to establish a suitable basis for the connection to short-distance (SD) contributions and lattice QCD Bruno et al. (2018), most conveniently expressed in terms of the chiral low-energy constants (LECs). For the real-emission diagrams, we critically revisit previous calculations using resonance models Flores-Baéz et al. (2006, 2007); Davier et al. (2010); Miranda and Roig (2020); Davier et al. (2024); Castro et al. (2025), both regarding the numerical implementation near threshold and the estimate of the resulting uncertainties, as described in Sec. 6. For the numerical evaluation, we then extend the dispersive representation of the pion form factor to include the $\rho^{\prime}$, $\rho^{\prime\prime}$ resonances, based on which we perform fits to the available data sets for the $\tau\to\pi\pi\nu_{\tau}$ spectral function from Belle Fujikawa et al. (2008), ALEPH Schael et al. (2005); Davier et al. (2014), CLEO Anderson et al. (2000), and OPAL Ackerstaff et al. (1999) in Sec. 7. The consequences for a $\tau$-based evaluation of the $2\pi$ contribution, $a_{\mu}^{\text{HVP, LO}}[\pi\pi,\tau]$, are discussed in Sec. 8, before concluding in Sec. 9. Further details on the radiative process $\tau\to\pi\pi\nu_{\tau}\gamma$ are relegated to App. A.
@@ 2 Formalism
Here and throughout we follow the convention that a pion without index is identified with a charged pion, in particular, $M_{\pi}\equiv M_{\pi^{\pm}}$ and thus $\beta_{\pi\pi}(s)=\sigma_{\pi}(s)$. The final factor in Eq. (12) refers to IB in the matrix elements, i.e., the pion VFF as measured in $e^{+}e^{-}$ scattering, $F_{\pi}^{V}(s)$, or the weak decay, $f_{+}(s)$. In the latter case, in principle, a second form factor $\tilde{f}_{-}(s,t)$ contributes to the matrix element, see Eq. (3.1) below, but only at second order in IB. At LO in IB, the $\tau$ decay rate can therefore be expressed as
@@
$\Delta(s,t)=1+2\,\text{Re}\,\tilde{f}_{+}(s,t)+g_{\text{rad}}(s,t)\,,\qquad g_{\text{rad}}(s,t)=g_{\text{Low}}(s,t)+g_{\text{rest}}(s,t)\,$, (18)
contains both virtual and real photon corrections, $\tilde{f}_{+}(s,t)$ and $g_{\text{rad}}(s,t)$, respectively (see below for their definition), where the latter is further separated into the leading contribution from Low’s theorem Low (1958) and the rest.
@@ 3.1 General considerations
$\mathcal{M}$ $=G_{F}V_{ud}^{*}f_{+}(s)\bar{u}(l_{2},\nu_{\tau})\gamma^{\mu}(1-\gamma_{5})u(l_{1},\tau)$
$\qquad\times\left[(q_{2}-q_{1})_{\mu}\big(1+\tilde{f}_{+}(s,t)\big)+(q_{1}+q_{2})_{\mu}\tilde{f}_{-}(s,t)\right]\,$, (23)
where $1+\tilde{f}_{+}(s,t)$ represents the contribution due to the $J^{P}=1^{-}$ component of the weak current and $\tilde{f}_{-}(s,t)$ vanishes in the isospin limit. We have factored out $f_{+}(s)$, which is defined in analogy to Eq. (6)
@@
normalized in such a way that $f_{+}(s)=F_{\pi}^{V}(s)$ in the isospin limit, and also separated the $\mathcal{O}(e^{2})$ corrections represented by $\tilde{f}_{+}(s,t)$. In these conventions, $\tilde{f}_{+}(s,t)\to 0$ in the limit in which the matrix element (3.1) is mediated by a single current insertion. In principle, Eq. (24) also involves a second form factor $f_{-}(s)$, which is related to $\tilde{f}_{-}(s,t)$ in the same single-current approximation and constitutes an example of a second-class current Weinberg (1958); Descotes-Genon and Moussallam (2014). As shown in Ref. Descotes-Genon and Moussallam (2014), $f_{-}(s)$ can be traded for a scalar form factor $f_{0}(s)$, with $f_{0}(0)=1$ and $f_{-}(s)=\frac{\Delta_{\pi}}{s}\big[f_{+}(s)-f_{0}(s)\big]$, which shows that $f_{-}(s)$ scales with the pion mass difference $\Delta_{\pi}=M_{\pi}^{2}-M_{\pi^{0}}^{2}$.
@@
$\tilde{f}_{+}^{\text{tree}}(s,t)=\tilde{f}_{-}^{\text{tree}}(s,t)=0\,,\qquad f_{+}^{\text{tree}}(s)=1\,.$ (26)
The spin-averaged squared amplitude can be expressed by
$\frac{1}{2}\sum\limits_{\text{spins}}|\mathcal{M}|^{2}$ $=2G_{F}^{2}|V_{ud}|^{2}|f_{+}(s)|^{2}\Big\{m_{\tau}^{2}(m_{\tau}^{2}-s)|\tilde{f}_{-}(s,t)|^{2}$ (27)
$-2m_{\tau}^{2}(2M_{\pi^{0}}^{2}+m_{\tau}^{2}-s-2t)\text{Re}\,\left[\big(1+\tilde{f}_{+}(s,t)\big)\tilde{f}_{-}^{*}(s,t)\right]$
$+[4M_{\pi}^{2}(M_{\pi^{0}}^{2}-t)+4t(s+t-M_{\pi^{0}}^{2})-m_{\tau}^{2}(s+4t)+m_{\tau}^{4}]\big|1+\tilde{f}_{+}(s,t)\big|^{2}\Big\}\,.$
Working at LO in IB, we use the formula above with $\tilde{f}_{-}(s,t)\to 0$ and $\big|1+\tilde{f}_{+}(s,t)\big|^{2}\to 1+2\,\text{Re}\,\tilde{f}_{+}(s,t)$ in the following. Furthermore, the differential decay width is given by
@@
with the phase space in the $\pi^{0}\nu_{\tau}$ invariant mass squared limited by $t_{\text{min}}(s)\leq t\leq t_{\text{max}}(s)$, see Eq. (16), and in the $\pi^{-}\pi^{0}$ invariant mass squared by $(M_{\pi}+M_{\pi^{0}})^{2}\leq s\leq m_{\tau}^{2}$. Reducing Eq. (3.1) to the single-current-insertion limit $f_{+}(s)\tilde{f}_{-}(s,t)\to f_{-}(s)$, $\tilde{f}_{+}(s,t)\to 0$, see Eq. (24), the integral over $t$ reproduces Cirigliano et al. (2001)
@@ 3.2 Radiative corrections
At $\mathcal{O}(e^{2}p^{2})$ in ChPT, the diagrams arising from the LO effective Lagrangian of Ref. Knecht et al. (2000) are shown in Fig. 1. In this work, we employ dimensional regularization to deal with both the UV and the IR divergences arising in the computation. Following the definition in Eq. (3.1), the form factors $\tilde{f}_{+}$ for the diagrams in Fig. 1 can be expressed as
@@ 4 Dispersive analysis
As a first modification beyond the pure ChPT analysis summarized in Sec. 3, it is clear that the strong final-state interactions between $\pi^{-}$ and $\pi^{0}$ need to be taken into account, after all, the main purpose of studying this decay concerns a precision measurement of the corresponding pion VFF. This leads to a nontrivial $f_{+}(s)$ in Eq. (3.1), while in the ChPT analysis of Sec. 3 one has $f_{+}(s)=1$ at this order in the expansion.
@@ 4.2 Dispersive calculation of the box diagram
When supplementing the box diagram with unsubtracted dispersive representations of the pion VFF, see Eq. (34), and projecting it onto the $\tilde{f}_{+}(s,t)$ form factor, see Eq. (3.1), the following expression can be obtained
@@ 4.3 Numerical treatment of $\boldsymbol{D_{0}}$ functions
The dependence on the (high-energy) integral cutoff $\Lambda^{2}$ of $I_{\ell 1}$ and the one in the integral $\int\text{d}s^{\prime\prime}D_{0}^{\text{rest}}(t,s^{\prime\prime})$ in Eq. (54) cancels against the $\Lambda$ dependence introduced via $\theta(\Lambda^{2}-s^{\prime\prime})$ in the second term proportional to $p_{2}(s,t)$. The imaginary part that emerges from the Cauchy propagator, see Eq. (55), also needs to be kept, as it contributes to $\text{Re}\,\tilde{f}_{+}(s,t)$ due to the imaginary part of $f_{+}(s)$ in Eq. (47).
@@ 7 Fits to the $\boldsymbol{\tau}$ spectral function
$\Delta(s,t)=1+2\,\text{Re}\,\tilde{f}_{+}^{\text{match}}(s,t)+g_{\text{Low}}(s,t)+g_{\text{rest}}(s,t)\,.$ (112)
@@ 9 Summary and conclusions
A further main outcome of this work is the substantial reduction of the uncertainty in the $G_{\text{EM}}(s)$ contribution, leaving the matching between the short-distance factor $S_{\text{EW}}^{\pi\pi}$ and the radiative corrections described by $G_{\text{EM}}(s)$ as the dominant source of uncertainty in the $\tau$-specific IB corrections. This matching can be further improved using input from lattice QCD, and establishing the latter connection could also help address the remaining, most critical IB correction in the matrix elements. That is, our work allows for a reliable calculation of the long-range radiative corrections, but lattice-QCD techniques as well as complementary dispersive calculations of IB in the pion VFF, see Ref. Colangelo et al. (2025b) for a first step, are also needed to achieve a complete account of IB corrections to hadronic $\tau$ decays. Once all these ingredients become available, a robust evaluation of the two-pion HVP contribution to the anomalous magnetic moment of the muon will be possible.